A gardener is installing a circular garden. The radius of the circle 28 feet. He needs to know the circumference in order to buy materials for the edging. What is the circumference if he uses = 22/7.
How much is the circumference?

Answer by mananth(16946)

Circumference of circle = 2*pi*r
r= 28
pi = 22/7
circumference = 2*28*28*22/7
=4,928 feet
..